# SmartPretrain: Model-Agnostic and Dataset-Agnostic Representation Learning for Motion Prediction

### v1

@@ Abstract
Predicting the future motion of surrounding agents is essential for autonomous vehicles (AVs) to operate safely in dynamic, human-robot-mixed environments. However, the scarcity of large-scale driving datasets has hindered the development of robust and generalizable motion prediction models, limiting their ability to capture complex interactions and road geometries. Inspired by recent advances in natural language processing (NLP) and computer vision (CV), self-supervised learning (SSL) has gained significant attention in the motion prediction community for learning rich and transferable scene representations. Nonetheless, existing pre-training methods for motion prediction have largely focused on specific model architectures and single dataset, limiting their scalability and generalizability. To address these challenges, we propose SmartPretrain, a general and scalable SSL framework for motion prediction that is both model-agnostic and dataset-agnostic. Our approach integrates contrastive and reconstructive SSL, leveraging the strengths of both generative and discriminative paradigms to effectively represent spatiotemporal evolution and interactions without imposing architectural constraints. Additionally, SmartPretrain employs a dataset-agnostic scenario sampling strategy that integrates multiple datasets, enhancing data volume, diversity, and robustness. Extensive experiments on multiple datasets demonstrate that SmartPretrain consistently improves the performance of state-of-the-art prediction models across datasets, data splits and main metrics. For instance, SmartPretrain significantly reduces the MissRate of Forecast-MAE by 10.6%. These results highlight SmartPretrain’s effectiveness as a unified, scalable solution for motion prediction, breaking free from the limitations of the small-data regime. Codes are available at our webpage.
@@ 1 Introduction
  • In the domain of generative SSL, pretext tasks are designed to learn context-rich scene representations from real data. Various traffic-related pretext tasks have been proposed, such as classification for maneuver or forecasting success  (Bhattacharyya et al., 2022), map graph relationship prediction (Azevedo et al., 2022), and traffic event detection (Pourkeshavarz et al., 2023). Moreover, extensive efforts have been made to replicate the success of Masked Autoencoder techniques seen in NLP and CV, exploring diverse traffic-specific masking strategies (Yang et al., 2023; Chen et al., 2023; Lan et al., 2024; Cheng et al., 2023). However, these pre-training strategies usually impose restrictions on the model architecture - they start by proposing pre-training strategies, whose realization then necessitates the model to generate certain types of features. For example, the detection for maneuvering, graph relationships, and traffic events require explicit formulation of these concepts/features within the model, and MAE approaches demand that each trajectory and map segment must have an explicit feature representation to enable reconstructive pre-training. Due to such inflexibility, these SSL strategies can usually only apply to specific models or, in most cases, their single specific model.
@@ 2.2 Self-Supervised Learning in Trajectory Prediction
Self-supervised learning (SSL) methods have been applied widely in both visual understanding (Chen et al., 2020; Grill et al., 2020; Caron et al., 2021; He et al., 2022; Benaim et al., 2020; Bertasius et al., 2021; Feichtenhofer et al., 2022) and natural language processing (Devlin et al., 2019; Brown, 2020; Touvron et al., 2023). SSL aims to learn informative and general representations via carefully designed pretext tasks, which can be fine-tuned for downstream tasks with supervision. Recently, the trajectory prediction community has made significant strides in incorporating SSL techniques. These methods utilize pretext tasks to pre-train models, allowing them to learn valuable representations that can be fine-tuned for enhanced trajectory prediction performance. Existing pre-training pipelines for trajectory prediction can be divided into three different categories, involving augmented or synthetic data (Yang et al., 2023; Li et al., 2024), contrastive learning (Xu et al., 2022; Bhattacharyya et al., 2022; Azevedo et al., 2022; Pourkeshavarz et al., 2023) and generative masked representation learning (Chen et al., 2023; Cheng et al., 2023; Lan et al., 2024).
@@ 3.2.1 Dataset-Agnostic Sampling
In the dataset-agnostic temporal sampling process, we aim to create positive and negative pairs for contrastive learning and construct trajectories for reconstructive learning, while ensuring data consistency across various datasets.
@@ 3.2.2 Model-Agnostic Contrastive and Reconstructive SSL
where $r$ denotes the cosine similarity measure, and $\tau$ denotes the temperature hyper-parameter. Specifically, the numerator term represents the similarity of positive sample pairs $(\mathbf{z}_{i,t},\mathbf{z}^{\prime}_{i,t^{\prime}})$, namely trajectory embeddings from the same agent in different timelines. The denominator term consists of two types of negative pairs: 1) intra-repelling pairs $(\mathbf{z}_{i,t},\mathbf{z}_{j,t})$: the trajectory embeddings from other agents of the same sub-scenario; 2) inter-repelling $(\mathbf{z}_{i,t},\mathbf{z}^{\prime}_{j,t^{\prime}})$: trajectory embeddings from different sub-scenarios. Through this objective, the similarities of the same agent’s embeddings are maximized, and those of the other pairs are minimized, thereby refining the model’s ability to capture meaningful contextual relationships and temporal dynamics.
@@ 3.2.3 Training Details
where $\lambda$ is a hyper-parameter balancing the contribution of both tasks.
@@ 4.1 Experimental Setup
Datasets. We train and evaluate our method on three large-scale motion forecasting datasets: Argoverse (Chang et al., 2019), Argoverse 2 (Wilson et al., 2023) and Waymo Open Motion Dataset (WOMD) (Sun et al., 2020). Argoverse contains 333k scenarios collected from interactive and dense traffic. Each scenario provides the HD map and 2 seconds of historic trajectory data, to predict the trajectory for the next 3 seconds, sampled at 10Hz. The training, validation, and test set of Argoverse is set to 205k, 39k and 78k scenarios, respectively. Argoverse 2 upgrades the previous dataset to include 250K sequences with higher prediction complexity. It extends the historic and prediction horizon to 5 seconds and 6 seconds respectively, sampled at 10Hz. The data is split into 200k, 25k, and 25k for training, validation, and test, respectively. For WOMD, the dataset provides 1 second of historic trajectory data and aims to predict the trajectory for 8 seconds into the future, sampled at 10Hz as well. It contains 487k training scenes, 44k validation scenes and 44k testing scenes. Notably, our pre-training pipeline only utilizes the training splits of these datasets.
@@
Implementation Details. We implement the projector and contrastive predictor as a 2-layer MLP with batch normalization, and the trajectory decoder also as a 2-layer MLP but with layer normalization to better fit its sequence nature. We use AdamW to optimize the online branch, which consists of the encoder in the motion prediction model, a projector, a contrastive predictor, and a trajectory decoder. In the momentum branch, the weights of the motion prediction encoder and the projector are initialized to be identical to that of the online branch, and updated via an exponential moving average (EMA) strategy. We use a momentum value of 0.996 and increase this value to 1.0 with a cosine schedule. We conduct pre-training with 32 Nvidia A100 40GB GPUs for 128 epochs, and finetuning with 8 Nvidia A100 40GB GPUs with the models’ original training schedules.
@@ 4.3 Ablation Studies
Figure 4: Visulization Results. The blue arrows are the model’s multi-modal trajectory predictions for the target agent, and the pink arrow is the ground truth future trajectory. After pre-training, the predicted trajectories get closer to the ground truth.
@@
Reconstrction Strategies. Our pre-training pipeline reconstructs the trajectories of the other sub-scenario $t^{\prime}$. For the ablation study, we explore more reconstruction targets, categorized into two groups. The first group incorporates predicting historical information, including the trajectory of the input sub-scenario and the entire scenario. The second group focuses on predictive reconstruction, excluding any historical data, and includes the complementary trajectory steps of the current sub-scenario and the trajectory of the other sub-scenario $t^{\prime}$. As shown in Table 5, pre-training with reconstruction tasks enhances prediction performance, with predictive targets yielding the most significant improvements.
Visulization. We present some visualization results in Fig. 4, which shows the prediction error of the target agent’s trajectories is consistently reduced after pre-training. With more datasets used in pre-training, the trajectories get closer to the ground truth, showing enhanced generalization and robustness with multiple datasets pre-training.

### v2

@@ Abstract
Predicting the future motion of surrounding agents is essential for autonomous vehicles (AVs) to operate safely in dynamic, human-robot-mixed environments. However, the scarcity of large-scale driving datasets has hindered the development of robust and generalizable motion prediction models, limiting their ability to capture complex interactions and road geometries. Inspired by recent advances in natural language processing (NLP) and computer vision (CV), self-supervised learning (SSL) has gained significant attention in the motion prediction community for learning rich and transferable scene representations. Nonetheless, existing pre-training methods for motion prediction have largely focused on specific model architectures and single dataset, limiting their scalability and generalizability. To address these challenges, we propose SmartPretrain, a general and scalable SSL framework for motion prediction that is both model-agnostic and dataset-agnostic. Our approach integrates contrastive and reconstructive SSL, leveraging the strengths of both generative and discriminative paradigms to effectively represent spatiotemporal evolution and interactions without imposing architectural constraints. Additionally, SmartPretrain employs a dataset-agnostic scenario sampling strategy that integrates multiple datasets, enhancing data volume, diversity, and robustness. Extensive experiments on multiple datasets demonstrate that SmartPretrain consistently improves the performance of state-of-the-art prediction models across datasets, data splits and main metrics. For instance, SmartPretrain significantly reduces the MissRate of Forecast-MAE by 10.6%. These results highlight SmartPretrain’s effectiveness as a unified, scalable solution for motion prediction, breaking free from the limitations of the small-data regime. The code is available at https://github.com/youngzhou1999/SmartPretrain.
@@ 1 Introduction
  • In the domain of generative SSL, pretext tasks are designed to learn context-rich scene representations from real data. Various traffic-related pretext tasks have been proposed, such as classification for maneuver or forecasting success  (Bhattacharyya et al., 2022), map graph relationship prediction (Azevedo et al., 2022), and traffic event detection (Pourkeshavarz et al., 2023). Moreover, extensive efforts have been made to replicate the success of Masked Autoencoder techniques seen in NLP and CV, exploring diverse traffic-specific masking strategies (Yang et al., 2023; Chen et al., 2023; Lan et al., 2024; Cheng et al., 2023). However, these pre-training strategies usually impose restrictions on the model architecture - they start by proposing pre-training strategies, whose realization then necessitates the model to generate certain types of features. For example, the detection for maneuvering, graph relationships, and traffic events require explicit formulation of these concepts/features within the model. MAE approaches demand that each trajectory or map segment must have an explicit feature representation to enable reconstructive pre-training. In approaches where only trajectories are masked and reconstructed (Yang et al., 2023), the method remains simple and flexible but essentially acts as a data augmentation technique. While many works focus on aggregating agent embeddings and provide explicit access to them, explicit map embeddings are not always available (Zhou et al., 2023; Tang et al., 2024). Thus the map reconstruction pre-training strategy in MAE could be less general. Due to such inflexibility, these SSL strategies can usually only apply to specific models or, in most cases, their single specific model.
@@ 2.2 Self-Supervised Learning in Trajectory Prediction
Self-supervised learning (SSL) methods have been applied widely in both visual understanding (Chen et al., 2020; Grill et al., 2020; Caron et al., 2021; He et al., 2022; Benaim et al., 2020; Bertasius et al., 2021; Feichtenhofer et al., 2022; Lin et al., 2023; 2024; Tang et al., 2025), natural language processing (Devlin et al., 2019; Brown, 2020; Touvron et al., 2023) and multimodal representation learning (Qu et al., 2025). SSL aims to learn informative and general representations via carefully designed pretext tasks, which can be fine-tuned for downstream tasks with supervision. Recently, the trajectory prediction community has made significant strides in incorporating SSL techniques. These methods utilize pretext tasks to pre-train models, allowing them to learn valuable representations that can be fine-tuned for enhanced trajectory prediction performance. Existing pre-training pipelines for trajectory prediction can be divided into three different categories, involving augmented or synthetic data (Yang et al., 2023; Li et al., 2024), contrastive learning (Xu et al., 2022; Bhattacharyya et al., 2022; Azevedo et al., 2022; Pourkeshavarz et al., 2023) and generative masked representation learning (Chen et al., 2023; Cheng et al., 2023; Lan et al., 2024).
@@ 3.2.1 Dataset-Agnostic Sampling
In the dataset-agnostic sampling process, we aim to create positive and negative pairs for contrastive learning and construct trajectories for reconstructive learning, while ensuring data consistency across various datasets.
@@ 3.2.2 Model-Agnostic Contrastive and Reconstructive SSL
where $r$ denotes the cosine similarity measure, and $\tau$ denotes the temperature hyper-parameter. Specifically, the numerator term represents the similarity of positive sample pairs $(\mathbf{z}_{i,t},\mathbf{z}^{\prime}_{i,t^{\prime}})$, namely trajectory embeddings from the same agent in different timelines. The denominator term consists of two types of negative pairs: 1) intra-repelling pairs $(\mathbf{z}_{i,t},\mathbf{z}_{j,t})$: the trajectory embeddings from other agents of the same sub-scenario; 2) inter-repelling pairs $(\mathbf{z}_{i,t},\mathbf{z}^{\prime}_{j,t^{\prime}})$: trajectory embeddings from different sub-scenarios. Through this objective, the similarities of the same agent’s embeddings are maximized, and those of the other pairs are minimized, thereby refining the model’s ability to capture meaningful contextual relationships and temporal dynamics.
@@ 3.2.3 Training Details
where $\lambda$ is a hyper-parameter balancing the contribution of both tasks and we set $\lambda$ as 1.0.
@@ 4.1 Experimental Setup
Datasets. We train and evaluate our method on three large-scale motion forecasting datasets: Argoverse (Chang et al., 2019), Argoverse 2 (Wilson et al., 2023) and Waymo Open Motion Dataset (WOMD) (Sun et al., 2020). Argoverse contains 333k scenarios collected from interactive and dense traffic. Each scenario provides the HD map and 2 seconds of historic trajectory data, to predict the trajectory for the next 3 seconds, sampled at 10Hz. The training, validation, and test set of Argoverse is set to 205k, 39k and 78k scenarios, respectively. Argoverse 2 extends the historic and prediction horizon to 5 seconds and 6 seconds respectively, sampled at 10Hz. The data is split into 200k, 25k, and 25k for training, validation, and test, respectively. For WOMD, the dataset provides 1 second of historic trajectory data and aims to predict the trajectory for 8 seconds into the future, sampled at 10Hz as well. It contains 487k training scenes, 44k validation scenes and 44k testing scenes. Notably, our pre-training pipeline only utilizes the training splits of these datasets.
@@
Implementation Details. We implement the projector and contrastive predictor as a 2-layer MLP with batch normalization, and the trajectory decoder also as a 2-layer MLP but with layer normalization to better fit its sequence nature. We use AdamW to optimize the online branch. In the momentum branch, the weights of the motion prediction encoder and the projector are initialized to be identical to that of the online branch, and updated via an exponential moving average (EMA) strategy. We use a momentum value of 0.996 and increase this value to 1.0 with a cosine schedule. We conduct single dataset pre-training with 8 Nvidia A100 40GB GPUs and data-scaled pertaining with 32 GPUs, both for 128 epochs. For fine-tuning, we use 8 GPUs with the models’ original training schedules.
@@ 4.2 Quantitative Results
4.3 Ablation Studies
To further demonstrate the effectiveness of utilizing pre-training datasets of our SmartPretrain, we also conduct an ablation on how to use the additional datasets in the pre-training phase. We explored two pre-training settings: 1) We pre-train with the standard motion prediction task on additional datasets and then finetuned on the downstream target dataset. 2) we directly train the downstream target dataset and additional datasets with the standard motion prediction task. The results and discussion are provided in Appendix A.1 due to limited space.
@@ 4.3 Ablation Studies
Reconstrction Strategies. Our pre-training pipeline reconstructs the trajectories of the other sub-scenario $t^{\prime}$ given the input sub-scenario $t$. For the ablation study, we explore more reconstruction targets, categorized into two groups. The first group incorporates predicting historical information, including 1) the trajectory of the input sub-scenario $t$ and 2) the entire scenario. The second group focuses on predictive reconstruction, excluding any historical data, and includes 3) the complementary trajectory steps of the input sub-scenario $t$ and 4) the trajectory of the other sub-scenario $t^{\prime}$. As shown in Table 5, pre-training with reconstruction tasks enhances prediction performance, with predictive targets yielding the most significant improvements.
Visulization. We present some visualization results of fine-tuning in Appendix A.2, categorized into four distinct scenarios: 1) trajectory alignment, 2) long trajectory prediction, 3) novel behavior generation, and 4) smooth and safe trajectory synthesis. These results demonstrate the model’s multi-modal trajectory predictions across various scenarios and highlight the enhanced generalization and robustness of the proposed method. We also show some reconstructed trajectories of pre-training in Appendix A.3 to illustrate the effectiveness of our pre-text task learning.
@@ 5 Conclusion
Appendix A Appendix
A.1 Ablation on how to pre-train additional datasets
We sincerely thank Reviewer TuBk for this contributive suggestion on our paper. The original discussion can be seen at: https://openreview.net/forum?id=Bmzv2Gch9v&noteId=xCpSNXeAyo.
We use Argoverse 2 as additional data to pre-train HiVT and Argoverse as the downstream target dataset, we explored two pre-training settings:
  • We pre-train the model on Argoverse 2 with the standard motion prediction task, and then fine-tune it to Argoverse.
  • We directly train the model on Argoverse and Argoverse 2 with the standard motion prediction task.
A minor design choice is that Argoverse 2 has a longer trajectory horizon than Argoverse. When pre-training on Argoverse 2, we could either randomly sample trajectory segments from the full trajectory, or use a fixed time window. For more comprehensive exploration, we explored both. For the fixed time window choice, considering Argoverse 2 data has 110 waypoints and Argoverse requires 50 waypoints (20 as inputs and 30 as outputs), we use Argoverse 2’s original current timestep, and collect 20 historic waypoints as input and 30 future waypoints as output.
We show the results from the first approach (random window) in Table 6 and the results from the second approach (fixed window) in Table 7.
Backbone Pre-training Dataset Finetuning Dataset Validation Set: Argoverse
minFDE $\downarrow$ minADE $\downarrow$ MR $\downarrow$
HiVT \ Argoverse 0.969 0.661 0.092
Argoverse 2 Argoverse 1.077 0.701 0.112
\ Argoverse + Argoverse 2 3.359 2.092 0.636
Table 6: Ablation study on motion prediction as the pre-text task for pre-training additional datasets (Argoverse 2 random window). Pre-training with the motion prediction task and then finetuning or directly training with mixed data does not boost prediction performance. They achieve worse prediction results than training from a single dataset.
Backbone Pre-training Dataset Finetuning Dataset Validation Set: Argoverse
minFDE $\downarrow$ minADE $\downarrow$ MR $\downarrow$
HiVT \ Argoverse 0.969 0.661 0.092
Argoverse 2 (fixed) Argoverse 1.078 0.697 0.112
\ Argoverse + Argoverse 2 (fixed) 1.214 0.762 0.133
Table 7: Ablation study on motion prediction as the pre-text task for pre-training additional datasets (Argoverse 2 fixed window). Pre-training with the motion prediction task and then finetuning or directly training with mixed data does not boost prediction performance. They achieve worse prediction results than training from a single dataset.
Interestingly, for both approaches, poor performance is observed when we directly use motion prediction as the pretraining task, or directly train the model from a mix of the two datasets (especially when we random sample from the additional dataset). It could be presumably due to: 1) the features learned by motion prediction are less transferable or robust, compared to the features learned from SSL tasks; 2) the trajectory distribution between different datasets is quite different, and could be pronounced when pre-training is performed on the standard prediction task.
A.2 Additional Visualization Results
We present additional visualization results to demonstrate the effectiveness of our method. These results are categorized into four distinct scenarios: (1) trajectory alignment, (2) long trajectory prediction, (3) novel behavior generation, and (4) smooth and safe trajectory synthesis.
Trajectory Alignment. As illustrated in Fig. 4, we showcase scenarios involving the prediction error of the target agent’s trajectories is consistently reduced after pre-training. With more datasets used in pre-training, the trajectories get closer to the ground truth, showing enhanced generalization and robustness with multiple datasets pre-training.
Figure 4: Visulization Results of Trajectory Alignment. The blue arrows are the model’s multi-modal trajectory predictions for the target agent, and the pink arrow is the ground truth future trajectory. After pre-training, the predicted trajectories get closer to the ground truth.
Long Trajectories. As illustrated in Fig. 5, we showcase scenarios involving long trajectory predictions compared against ground truth data. Incorporating additional datasets during pre-training enhances the model’s ability to capture fine-grained spatiotemporal interactions with objects within the scenario, thereby enabling more accurate trajectory predictions in the long term.
Figure 5: Visulization Results of Long Trajectories. The blue arrows are the model’s multi-modal trajectory predictions for the target agent, and the pink arrow is the ground truth future trajectory. Our pre-training method enables more accurate trajectory predictions in the long term.
Behavior Generation. We present some scenarios in which new behavior is learned through pre-training. In Fig. 6, the original model struggles to align well with the ground truth across six modality outputs. However, after pre-training, particularly multiple datasets pre-training, the model acquires new behaviors, resulting in the generation of more meaningful and skillful trajectories. For example, lane-changing maneuvers (Fig. 6 top), lane-keeping (Fig. 6 middle) and trajectories that are closely synchronized with the ground truth (Fig. 6 bottom).
Figure 6: Visulization Results of Behavior generation. The blue arrows are the model’s multi-modal trajectory predictions for the target agent, and the pink arrow is the ground truth future trajectory. Our pre-training method can teach new behavior learned through multiple datasets to the model, thus generating more meaningful and skillful trajectories.
Smoothness and Safety Synthesis. In Fig. 7, we illustrate scenarios where pre-training facilitates the generation of smoother and safer trajectories compared to models without pre-training. Coarse trajectories are improved into smoother and safer ones, demonstrating the generalizability and robustness of the learned features achieved through our multiple datasets pre-training.
Figure 7: Visulization Results of Smooth and Safety Synthesis. The blue arrows are the model’s multi-modal trajectory predictions for the target agent, and the pink arrow is the ground truth future trajectory. Our pre-training method contributes to learning smoother and safer trajectories, especially through multiple datasets pre-training.
A.3 Reconstructed Trajectories
We present visualization results of reconstructed trajectories to demonstrate the effectiveness of our pre-text task learning. As shown in Fig. 8, the reconstructed trajectories align closely with the ground truth, indicating the successful learning of our pretext task.
Figure 8: Visulization Results of Reconstructed Trajectories. Left: The blue arrows are the trajectories of sub-scenario $t$, and the pink arrows are the trajectories of the entire scenario. Right: The blue arrows are the reconstructed trajectories of sub-scenario $t^{\prime}$, and the pink arrows are the ground truth trajectories of the sub-scenario $t^{\prime}$. Through pre-training, the model successfully reconstructs the sub-scenario $t^{\prime}$ based on sub-scenario $t$. The promising results suggest that the reconstruction task has been effectively learned.